Calculate the electric field at one corner off a square 80π‘π‘š on a side if the other three corners are occupied by charges each of magnitude 18.2 x 10-4C.


1
Expert's answer
2021-06-15T15:11:26-0400

Side of square, a = "80\space cm=80\times10^{-2}\space m"

Electric field due to horizontally placed charge, "E_1=\dfrac{kq}{a^2}=\dfrac{(9\times10^9)(18.2\times10^{-4})}{(80\times10^{-2})^2}"

"E_1=2.0475\times10^{9}\space N/C(left)"


Electric field due to vertically placed charge, "E_2=2.0475\times10^{9}\space N/C(downward)"

"E_{res_1}=\sqrt{E_1^2+E_2^2+2E_1E_2\cos90\degree}=3.50\times10^9\space N/C(diagonally)"


Electric field due to diagonally placed charge, "E_3=\dfrac{(9\times10^9)(18.2\times10^{-4})}{(80\sqrt{2}\times10^{-2})^2}"

"E_3=1.447\times10^{9}\space N/C"


"E_{res}=E_{res_1}+E_3=4.94\times10^{9}\space N/C(diagonally)"